OpenStudy (anonymous):

Solve this system of equation. y = −x2 + 4 y = 2x + 1

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=-x^2+4 }\) \(\large\color{black}{ \displaystyle y=2x+1 }\) We know that both \(\large\color{black}{ \displaystyle 2x+1 }\) and \(\large\color{black}{ \displaystyle -x^2+4 }\) are equivalent to \(\large\color{black}{ \displaystyle y }\), therefore we can set up the following: \(\large\color{black}{ \displaystyle 2x+1=-x^2+4 }\)

OpenStudy (solomonzelman):

rearrange this, \(\large\color{black}{ \displaystyle x^2+2x-3=0 }\)

OpenStudy (solomonzelman):

Now, you just need to solve this quadratic for x, and then solve for y.